what will be the length of the perpendicular drawn from the point (-3,-4) to the straight line 12(x+6) = 5(y-2)

Grade:12th pass

2 Answers

Arun
6 years ago
for the length of the perpendicular
Waseem khan
6 years ago
LET US FIRST SOLVE THE EQUATION OF LINE THAT IS 12X-5Y+82=0 LET THE POINT X1=-3ANDY1=-4THE LENGTH OF THE PERPENDICULAR WILL BE 12X1-5Y1+82/√(12²+5²)12*-3-5*-4+82/13=66/13So the lenght of perpendicular is 66/13=5.07